# LocMoE+: Enhanced Router with Token Feature Awareness for Efficient LLM Pre-Training

### v1

@@ Abstract
Mixture-of-Experts (MoE) architectures have recently gained increasing popularity within the domain of large language models (LLMs) due to their ability to significantly reduce training and inference overhead. However, MoE architectures face challenges, such as significant disparities in the number of tokens assigned to each expert and a tendency toward homogenization among experts, which adversely affects the model’s semantic generation capabilities. In this paper, we introduce LocMoE+, a refined version of the low-overhead LocMoE, incorporating the following enhancements: (1) Quantification and definition of the affinity between experts and tokens. (2) Implementation of a global-level adaptive routing strategy to rearrange tokens based on their affinity scores. (3) Reestimation of the lower bound for expert capacity, which has been shown to progressively decrease as the token feature distribution evolves. Experimental results demonstrate that, without compromising model convergence or efficacy, the number of tokens each expert processes can be reduced by over 60%. Combined with communication optimizations, this leads to an average improvement in training efficiency ranging from 5.4% to 46.6%. After fine-tuning, LocMoE+ exhibits a performance improvement of 9.7% to 14.1% across the GDAD, C-Eval, and TeleQnA datasets.
1 Introduction
@@
Large language models (LLMs) have shown exceptional proficiency in understanding deep structures and complex semantic relationships within language [1]. As these models scale up, their capabilities in language generation and logical comprehension are enhanced, but this comes at the cost of significant computational, communication, and storage demands [2]. To scale models efficiently without disproportionately increasing computational costs, researchers have incorporated the Mixture-of-Experts (MoE) architecture into LLMs [3]. The MoE framework integrates multiple experts within the model, each tasked with processing specific types of inputs [4]. For a given input, only a subset of experts is activated, allowing for more efficient use of computational resources [5]. Recently, several LLMs employing MoE structures, such as DeepSeek-MoE [6] and Mixtral [7], have demonstrated outstanding performance on various leaderboards.
Despite the efficiency benefits of MoE in scaling model sizes, it introduces several new challenges and drawbacks [8]. For instance, an imbalanced "winner-takes-all" phenomenon among experts can detrimentally affect overall model performance [9], and the All-to-All communication pattern can become a bottleneck [10]. The model’s convergence and the experts’ generalization capabilities are heavily dependent on the design of the routing strategy. Recent research has addressed these challenges from multiple angles, such as optimizing routing strategies and enhancing communication efficiency. In terms of routing strategies, DeepSeek-MoE [6] introduced the concepts of vertical-class experts, created through fine-grained partitioning, and shared experts, which are always activated. Google Brain proposed the Expert Choice (EC) routing algorithm [11], assigning experts with predetermined buffer capacities to the Top-k tokens to ensure load balance. For communication efficiency, FastMoE [12] leverages Megatron-LM [13], a distributed training framework by NVIDIA, utilizing CUDA’s multi-stream mechanism to maximize the overlap between communication and computation. Tutel [14] introduced flexible All-to-All communication to alleviate throughput bottlenecks caused by smaller matrix dimensions and proposed 2DH All-to-All to align non-contiguous memory spaces and merge scattered communications. The uneven expert load in classical MoE [15] often necessitates token dropping or padding. Megablocks [16] employs block-sparse operations to overcome the limitations of dynamic routing, improving the balance between model quality and efficiency.
LocMoE [17] aims to enhance the interpretability of routing mechanisms while reducing computational and communication overhead. LocMoE leverages orthogonal routing weights to prevent token homogenization across different expert networks and introduces the Grouped Average Pooling (GrAP) layer [18] for token feature extraction. Under these conditions, LocMoE uses theoretical proofs to calculate the lower bound of expert capacity. Experimental results demonstrate that employing this lower bound does not degrade, but slightly enhances, performance on downstream tasks. Additionally, LocMoE introduces a loss regularization term that incorporates data locality and load balancing [19], leading to a 12% to 22% improvement in training efficiency.
Building on previous work, this paper proposes LocMoE+, which aims to further enhance training efficiency and model performance. LocMoE+ adopts GrAP and the locality loss regularization terms from LocMoE, and continues to analyze the impact of token features on expert selection. The primary contributions of this paper are as follows:
@@
    The affinity score. Given that expert selection is related to the angle between tokens and gating weights, the cosine similarity between tokens and gating weights is used to define the affinity score between tokens and experts. This score is utilized for token rearrangement and as a threshold for the lower bound of expert capacity.
@@
    The hybrid token rearrangement scheme. By integrating the concepts of Token Choice Router (TCR) and Expert Choice Router (ECR), the training success rate is enhanced while considering expert capacity constraints. Its effectiveness has been theoretically validated.
@@
    The adaptive expert capacity. Setting a holistic affinity threshold allows the lower bound of expert capacity to be significantly reduced. As training iterations increase, the information density of token features grows, causing the expert capacity to initially decrease and then stabilize, thereby enhancing training efficiency.
LocMoE+ uses the state-of-the-art MoE model Mixtral 8$\times$7B as the backbone, and utlizes AscendSpeed and ModelLink libraries for pre-training on Ascend NPU clusters. Experiments conducted on clusters with 32, 64, and 256 NPUs show that LocMoE+ improves training efficiency by 5.4% to 46.6% compared to the baseline, and by 2.9% to 13.3% compared to LocMoE. Model performance is enhanced by 9.7% to 14.1% compared to the baseline, and by 1.7% to 4.1% compared to LocMoE.
The rest of this paper is structured as follows: Section 2 introduces related work on LLMs, MoE, and the Ascend architecture. Section 3 reviews LocMoE and presents the methods proposed in this paper, along with theoretical evidence. Section 4 analyzes the experimental results of LocMoE+ regarding training efficiency and model performance. The final section summarizes the content of this paper and offers an outlook on future improvements.
2 Related Work
@@
BERT [20] and GPT [21] in 2018 marked the emergence of the LLM era. GPT-3 [22], with 175 billion parameters, demonstrated impressive capabilities in language understanding and generation in 2020. By 2023-2024, LLMs expanded to hundreds of billions to trillions of parameters: GPT-4 [23] achieved breakthroughs in zero-shot and few-shot learning; PaLM-2 [24] enhanced knowledge acquisition and reasoning; Chinchilla [25] optimized computational resources and model scale; and Claude-v1 [26] displayed a safer, human-aligned LLM. These LLMs have made significant advancements across domains, paving the way towards artificial general intelligence.
@@
Despite the outstanding performance of LLMs in language understanding and generation, their high computational cost hinders further model scale expansion. To alleviate this issue, researchers have introduced the MoE architecture, which incorporates multiple expert modules and dynamically selects a subset of experts to process inputs, significantly reducing computational overhead while maintaining performance. GShard [3] and Switch Transformer [4] employ Top-2 and Top-1 routing strategies, respectively, to mitigate the expert load imbalance problem. Hash Layer [27] adopt hashing techniques to simplify the gating mechanism and improve model performance. METRO [28] introduces a metric learning-based routing mechanism that guides expert selection by learning task-relevant metric functions, enhancing the model’s generalization ability. DSelect-k [29] proposes a dynamic selection algorithm based on Markov chains, adaptively choosing the number of experts to maintain performance while reducing training costs. Furthermore, DeepSpeed-MoE [30] optimizes expert parallelism and pipeline parallelism in MoE, further boosting training efficiency. These works lay the foundation for the subsequent application of MoE in LLMs.
@@
The AI Core in the Ascend 910B NPU adopts the DaVinci architecture [31], which is responsible for executing computation-intensive operators related to scalars, vectors, and 3D cubes, and is widely used in LLM training. Huawei’s proprietary cache coherence protocol interconnect, High Confidence Computing Systems (HCCS) [32], enables high-performance inter-device data communication in multi-card scenarios. The Huawei Collective Communication Library (HCCL) [33], designed for the Ascend processor, offers high-performance primitives for single-node multi-card and multi-node multi-card collective communication. It implements collective communication functions on PCIe, HCCS, and RoCE high-speed links, facilitating distributed training. PyTorch [34], a widely adopted deep learning framework, is compatible with the Ascend NPU through the Ascend extension for PyTorch [35]. AscendSpeed, an acceleration library tailored for Ascend devices, focuses on parallel strategy partitioning for LLMs and aggregates results through collective communication. ModelLink, a training framework designed for LLMs, fully considers the characteristics of the Ascend NPU and implements features such as memory optimization, multi-dimensional parallelism, and adaptive data loading, further enhancing training performance.
3 Method
Notation.
[…]
3.1 Background
Main Principle.
The MoE architecture, based on the Transformer framework, efficiently scales up model size with low computational overhead, benefiting from two primary structures: a sparse gating network for routing tokens and expert networks for processing specific token categories. The sparse gating network determines the expert to which a token will be sent.
We consider the supervised classification for brevity where the training samples are $\{({\bm{x}}^{(i)},y_{i})\}_{i=1}^{N}\sim{\mathcal{D}}$. Each training sample ${\bm{x}}^{\top}=({\bm{x}}_{1}^{\top},\dots,{\bm{x}}_{s}^{\top})\in{\mathbb{R}}^{sd}$ has $s$ tokens with token feature ${\bm{x}}_{i}\in{\mathbb{R}}^{d},\forall i\in[s]$, and label $y\in{\mathbb{N}}^{+}$. The objective is to learn the map of ${\bm{x}}$ to the corresponding $y$. The general MoE structure are formulated as
@@ 3.1 Background
where $n$ is the number of experts, $G({\bm{x}}_{t})\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{n}$ is the gating (or router) weights of experts which maps the tokens of ${\bm{x}}_{t}$ into the coresponding experts with weights, e.g., $G_{i}({\bm{x}})=\text{Softmax}({\bm{W}}{\bm{x}}+\bm{\epsilon})$ where the softmax is applied to each row, and $E_{i}({\bm{x}}_{t})\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$ is the $i$-th expert network, e.g., simple MLP models, see [36] for current different router methods. Generally, $n\ll s$, which saves much computation compared to the dense structure.
@@
There are two main router methods: hard router [36] which uses hard assignments between tokens and experts with sparse dispatch matrix, and soft router [37] which uses weighted combinations of tokens. In this paper, we still focus on the first one, since the computation cost is much smaller than the soft ones.
The hard router also have two main choices: 1) Token Choice Router (TCR) [38, 11], which lets each token choose its top-scored experts; Taking Top1 routing as an example, it typically selects a linear layer to extract feature representations for different experts and then obtains the expert index with the highest routing probability through the gating function. 2) Expert Choice Router (ECR), which lets each expert choose its top-scored tokens. Moreover, to balance the budget of differnet experts, we will pre-define the expert capacity $C$ of all experts in practice, namely, the maximum tokens sent to each expert.
Chocie of score. In this paper, we define the score of $t$-th token and $i$-th expert as $\delta_{ti}=\cos\left({\bm{x}}_{t},{\bm{w}}_{i}\right):={\bm{x}}_{t}^{\top}{\bm{w}}_{i}/(\|{\bm{x}}_{t}\|\cdot\|{\bm{w}}_{i}\|)$, i.e., the cosine simlarity between vectors ${\bm{x}}_{t}$ and ${\bm{w}}_{i}$. The score is to decide the gating weight in router. For example, we can route $t$-th token to the experts of the $-\ell$ indices in $(\delta_{t1},\dots,\delta_{tn})$.
Chocie of capacity. In previous works, expert capacity $C$ is calculated using a capacity factor, ensuring that the size of each expert’s token buffer remains consistent. The choice of capacity factor is currently based mainly on empirical knowledge; increasing the capacity factor allows for handling more tokens but reduces training efficiency.
TCR expression. First, given $\ell$, we decide the candidate expert set for each token by the sort of scores:
$\left(\tilde{I}_{t1},\dots,\tilde{I}_{t\ell}\right)=\text{Top-}\ell\left(\left\{\delta_{t1},\dots,\delta_{tn}\right\}\right),\tilde{I}_{tk}\in[n],\forall t\in[s],k\in[\ell].$ (2)
Second, we choose the first-$C$ tokens (or all if less than $C$) to each expert due to the capacity limit:
$\left(I_{1i},\dots,I_{Ci}\right)=\text{Bottom-}C\left(\left\{t\in[s]:\exists j\in[\ell],\tilde{I}_{tj}=i\right\}\right),I_{ki}\in[s]\cup\text{None},\forall i\in[n],k\in[C].$ (3)
In practice, we adopt $\ell=1$ to further save the computation cost since $n\ll s$.
ECR expression. We directly choose the tokens in the top-$C$ scores of each expert:
$\left(I_{1i},\dots,I_{Ci}\right)=\text{Top-}C\left(\left\{\delta_{1i},\dots,\delta_{si}\right\}\right),I_{ki}\in[s],\forall i\in[n],k\in[C].$ (4)
Intuitively, TCR can result in underused experts with fewer tokens than their allowed capacity, while ECR can result in underused tokens where no expert receives such token.
Auxiliary Loss.
@@
Feed-forward network (FFN) layers are commonly employed in expert networks, allowing each expert to learn independently as a separate neural network, thus preventing interference between samples. This mechanism leads to a severe load imbalance, as experts frequently selected in the early stages are more likely to be chosen in later stages. To mitigate this skewness in token allocation, the auxiliary loss [8] has been proposed, defined as $L_{\mathrm{aux}}=\alpha{{n}{\sum_{i=1}^{n}f_{i}{P_{i}}}}$, where $f_{i}=\frac{1}{T}\sum_{{\bm{x}}\in{\beta}}\mathds{1}{\mathop{\arg\max}p({\bm{x}})=i}$, and $P_{i}=\frac{1}{T}\sum{{\bm{x}}\in{\beta}}p_{i}({\bm{x}})$. Here, $\beta$ represents the batch containing $T$ tokens, and $p_{i}$ denotes the probability of a token being routed to expert $i$.
3.2 LocMoE
GrAP.
[…]
The Lower Bound of the Expert Capacity.
@@ 3.2 LocMoE
Locality Loss.
Building upon the auxiliary loss, LocMoE introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$. This loss term serves as a soft constraint, encouraging tokens to be sent to experts residing on the same node, thereby mitigating the substantial overhead incurred by partial inter-node communication.
3.3 Token Awareness Router for LocMoE
3.3.1 Scheme Sketch
Figure 1 illustrates the model constructed in this paper, with a particular focus on the proposed token awareness router. The model framework is GPT-like, comprising only decoders, and incorporates LocMoE+ in place of the gated routing of sparse MoE blocks. To enhance interpretability and reduce computational overhead, the original Top-2 routing is simplified to the sparsest Top-1 routing. Consequently, the weights of GrAP, consisting of orthogonal vectors, emphasize different aspects of token features, which are evenly divided and non-overlapping. This routing mechanism can be viewed as a multi-classification of these feature fragments.
@@ 3.3.1 Scheme Sketch
TCR, introduced in the background, is exemplified in the figure. If the feature fragment corresponding to the $k$-th dimension of the gating weight for a particular token is more prominent, then that token will be routed to the $k$-th expert. If among all tokens routed to the $k$-th expert, there is a certain probability of the presence of class-discriminative tokens, then the capacity $C$ must be set to a larger value to ensure the inclusion of sufficient class-discriminative tokens. The router proposed in this paper is a hybrid of TCR and ECR modes. After determining the expert to which a token will be routed, scores are calculated for the tokens assigned to each expert, and a $-\ell$ selection is performed, where $\ell^{*}$ is determined by a threshold of the sum of scores. Subsequent theoretical analysis will demonstrate the effectiveness of this hybrid routing scheme.
Figure 1: The architecture of the model along with the hybrid TCR+ECR router.
3.3.2 Theoretical Explanation
@@
To explain the motivation of our method, we show some theoretical insights in this section. Our theoretical analyis is bulit on [39], where they assume the training data has unique class-discriminative patterns with exectly the same number as experts, and other class-irrelevant patterns. Chowdhury et al. [39] demonstrated that the classical MoE structure go through two phases:
@@
Phase 1: Router training ([39, Lemma 4.1, Assumption 4.4]), which makes class-discriminative patterns all to the corresponding expert. This process ensures that each expert only receives the class-discriminative tokens related to the specific class.
@@
Phase 2: Expert training ([39, Theorem 4.2, Theorme 4.5]), which makes each expert learn to predict the label based on its class-discriminative inputs from Phase 1. This process is designed to establish each expert’s ability to handle and solve problems.

### v2

# Expert-Token Resonance: Redefining MoE Routing through Affinity-Driven Active Selection

@@ Abstract
Mixture-of-Experts (MoE) architectures have emerged as a paradigm-shifting approach for large language models (LLMs), offering unprecedented computational efficiency. However, these architectures grapple with challenges of token distribution imbalance and expert homogenization, impeding optimal semantic generalization. We introduce a novel framework that redefines MoE routing through affinity-driven active selection. The innovations for the framework encompass: (1) A rigorous formulation of expert-token affinity metrics. (2) An adaptive bidirectional selection mechanism leveraging resonance between experts and tokens. (3) Theoretical derivation and experimental evidence of reduced expert capacity bounds under dynamic token distribution evolution. It is also integrated with orthogonal feature extraction module and an optimized loss function for expert localization. Our theoretical analysis demonstrates that this approach mitigates expert homogenization while enabling substantial capacity boundary reduction. Experimental validation corroborates these findings: it achieves a 40% reduction in token processed by each expert without compromising model convergence or efficacy. When coupled with communication optimizations, the training efficiency improvements of 5.4% to 46.6% can be observed. After supervised fine-tuning, it exhibits performance gains of 9.7% to 14.1% across GDAD, C-Eval, and TeleQnA benchmarks.
Introduction
@@
Large language models (LLMs) have shown exceptional proficiency in understanding deep structures and complex semantic relationships within language (Zhao et al. 2023). As these models scale up, their capabilities in language generation and logical comprehension are enhanced, but this comes at the cost of significant computational, communication, and storage demands (Jiang et al. 2024b). To scale models efficiently without disproportionately increasing computational costs, researchers have incorporated the Mixture-of-Experts (MoE) architecture into LLMs (Lepikhin et al. 2020). The MoE framework integrates multiple experts within the model, each tasked with processing specific types of inputs (Fedus, Zoph, and Shazeer 2022). For a given input, only a subset of experts is activated, allowing for more efficient use of computational resources (Du et al. 2022). Recently, several LLMs employing MoE structures, such as DeepSeek-MoE (Dai et al. 2024) and Mixtral (Jiang et al. 2024a), have demonstrated outstanding performance on various leaderboards.
Despite the efficiency benefits of MoE in scaling model sizes, it introduces several new challenges and drawbacks (Shazeer et al. 2017). The conventional MoE model’s convergence and the experts’ generalization capabilities are heavily dependent on the design of the routing strategy, which easily leads to an imbalanced ”winner-takes-all” phenomenon among experts. The imbalance between excessively ”developed” experts and those lacking adequate training may compromise or even nullify the intended functionality of routing strategies. Recent studies address these challenges from multiple perspectives (Li et al. 2023). DeepSeek-MoE (Dai et al. 2024) introduces the concept of the vertical-class expert, which is created through fine-grained partitioning, and the shared experts are always activated. This approach minimizes the constraints on expert specialization imposed by knowledge hybridity and knowledge redundancy. StableMoE (Dai et al. 2022) proposes a two-stage training approach to address the issue of routing fluctuation. This method involves training the routing network independently from the backbone model and utilizing a frozen, distilled routing mechanism to allocate tokens. The characteristics of classical gated routing lead to experts being unable to learn features mastered by other experts. To address this, MoDE (Xie et al. 2024) proposes moderate distillation between experts to mitigate the generalization problems caused by narrow learning paths. DYNMoE (Guo et al. 2024) introduces a unique gated routing mechanism capable of adaptively determining the number of activated experts through trainable expert thresholds, even allowing for the addition or removal of experts.
In addition to the classical token choice scenario, previous researches also propose work utilizing expert choice (EC). Google Brain introduces the EC routing algorithm (Zhou et al. 2022), which assigns experts with predetermined buffer capacities to the Top-k tokens to ensure load balance. The Brainformer (Zhou et al. 2023) also adopts this routing strategy, constructing a trainable gating matrix to project the input feature space onto scores corresponding to each expert. Then, each token is routed to the Top-k experts. This strategy is proven highly effective in achieving expert load balancing and enhancing expert learning outcomes.
The design of routing strategy is crucial to the MoE structure, while not all tokens may be suitable for training (Riquelme et al. 2021). In addition to data preprocessing techniques such as dataset cleaning and deduplication, previous studies have also considered how to discard certain tokens within the model. Early work introduced the concept of expert capacity (Lepikhin et al. ), which refers to the maximum number of tokens each expert can process at once. Tokens exceeding this capacity are discarded. Expert capacity helps to ensure load balance among experts while facilitating All-to-All communication implementation. However, in situations where it is uncertain whether a token contributes to training, there is a risk of discarding class-discriminative samples, potentially compromising the model’s training outcomes. XMoE (Yang et al. 2024) achieves more precise router by implementing a threshold-based approach. If a token reaches the specified threshold, it is processed exclusively by a single expert while being discarded by other experts within the Top-k selection. This method allows for more nuanced token selection and processing. LocMoE (Li et al. 2024) leverages orthogonal routing weights to prevent token homogenization across different expert networks and introduces the Grouped Average Pooling (GrAP) layer (Wang, Zhang, and Du 2023) for token feature extraction. Under these conditions, LocMoE also provides the theoretical proof for the lower bound of expert capacity.
In this paper, we propose expert-token resonance, a mechanism consisting of an expert-token bidirectional selection router and the adaptive expert capacity strategy. The primary contributions of this paper are as follows:
@@
    A rigorous formulation of expert-token affinity metrics. Given that expert selection is related to the angle between tokens and gating weights, the cosine similarity between tokens and gating weights is used to define the affinity score between tokens and experts. We provide theoretical analysis of utilizing the affinity score for better guiding expert choice routers to focus on different segmentations of tokens, thus mitigating the homogenization tendency. Further, we incorporate affinity scores into the design of our bidirectional selection router.
@@
    Expert-token bidirectional selection. By integrating the concepts of expert choice router (ECR) and token choice router (TCR), we propose the adaptive bidirectional selection mechanism. Contrast to conventional router, the bidirectional selection router allows MoE to enhance the training success rate while considering expert capacity constraints. Its effectiveness has been theoretically validated.
@@
    The adaptive expert capacity bound. Setting an adaptive affinity threshold allows the lower bound of expert capacity to be significantly reduced. As training iterations increase, the information density of token features grows, causing the expert capacity to initially decrease and then stabilize. Ultimately, the training efficiency of MoE can be greatly enhanced.
Expert-token resonance mechanism adopts the state-of-the-art MoE model Mixtral 8$\times$7B as the backbone, and utilizes AscendSpeed and ModelLink libraries for pre-training on Ascend NPU clusters. Experiments conducted on clusters with 32, 64, and 256 NPUs indicate that our approach improves training efficiency by 5.4% to 46.6% compared to the baseline, and by 2.9% to 13.3% compared to LocMoE. Model performance is enhanced by 9.7% to 14.1% compared to the baseline, and by 1.7% to 4.1% compared to LocMoE.
The rest of this paper is structured as follows: Section Related Work introduces related work on LLMs, MoE, and the Ascend architecture. Section Method reviews LocMoE and presents the methods proposed in this paper, along with theoretical evidence. Section Experiments analyzes the experimental results of our approach regarding training efficiency and model performance. The final section summarizes the content of this paper and offers an outlook on future improvements.
Related Work
@@
BERT (Kenton and Toutanova 2019) and GPT (Radford et al. ) in 2018 marked the emergence of the LLM era. GPT-3 (Floridi and Chiriatti 2020), with 175 billion parameters, demonstrated impressive capabilities in language understanding and generation in 2020. By 2023-2024, LLMs expanded to hundreds of billions to trillions of parameters: GPT-4 (Achiam et al. 2023) achieved breakthroughs in zero-shot and few-shot learning; PaLM-2 (Anil et al. 2023) enhanced knowledge acquisition and reasoning; Chinchilla (Hoffmann et al. 2022) optimized computational resources and model scale; and Claude-v1 (Wei, Haghtalab, and Steinhardt 2024) displayed a safer, human-aligned LLM. These LLMs have made significant advancements across domains, paving the way towards artificial general intelligence.
@@
Despite the outstanding performance of LLMs in language understanding and generation, their high computational cost hinders further model scale expansion. To alleviate this issue, researchers have introduced the MoE architecture, which incorporates multiple expert modules and dynamically selects a subset of experts to process inputs, significantly reducing computational overhead while maintaining performance. GShard (Lepikhin et al. 2020) and Switch Transformer (Fedus, Zoph, and Shazeer 2022) employ Top-2 and Top-1 routing strategies, respectively, to mitigate the expert load imbalance problem. Hash Layer (Roller et al. 2021) adopt hashing techniques to simplify the gating mechanism and improve model performance. METRO (Bajaj et al. 2022) introduces a metric learning-based routing mechanism that guides expert selection by learning task-relevant metric functions, enhancing the model’s generalization ability. DSelect-k (Hazimeh et al. 2021) proposes a dynamic selection algorithm based on Markov chains, adaptively choosing the number of experts to maintain performance while reducing training costs. Furthermore, DeepSpeed-MoE (Rajbhandari et al. 2022) optimizes expert parallelism and pipeline parallelism in MoE, further boosting training efficiency. These works lay the foundation for the subsequent application of MoE in LLMs.
@@
The AI Core in the Ascend 910B NPU adopts the DaVinci architecture (Tang and Wang 2023), which is responsible for executing computation-intensive operators related to scalars, vectors, and 3D cubes, and is widely used in LLM training. Huawei’s proprietary cache coherence protocol interconnect, High Confidence Computing Systems (HCCS) (Xia et al. 2021), enables high-performance inter-device data communication in multi-card scenarios. The Huawei Collective Communication Library (HCCL) (Cao et al. 2024), designed for the Ascend processor, offers high-performance primitives for single-node multi-card and multi-node multi-card collective communication. It implements collective communication functions on PCIe, HCCS, and RoCE high-speed links, facilitating distributed training. PyTorch (Paszke et al. 2019), a widely adopted deep learning framework, is compatible with the Ascend NPU through the Ascend extension for PyTorch (Zhu et al. 2023). AscendSpeed, an acceleration library tailored for Ascend devices, focuses on parallel strategy partitioning for LLMs and aggregates results through collective communication. ModelLink, a training framework designed for LLMs, fully considers the characteristics of the Ascend NPU and implements features such as memory optimization, multi-dimensional parallelism, and adaptive data loading, further enhancing training performance.
Method
In this section, we present the definition of affinity score with a rigorous formulation of expert-token affinity metrics and the implementation of our affinity-driven active selection routing. To reduce the lower bound of expert capacity, we introduce a training strategy for dynamic token distribution evolution under the guiding principle of affinity score.
Model Architecture.
Backbone.
The MoE architecture, based on the Transformer framework, efficiently scales up model size with low computational overhead, benefiting from two primary structures: a sparse gating network for routing tokens and expert networks for processing specific token categories.
We consider the supervised classification for brevity where the training samples are $\{(\bm{x}^{(i)},y_{i})\}_{i=1}^{N}\sim\mathcal{D}$. Each training sample $\bm{x}^{\top}=(\bm{x}_{1}^{\top},\dots,\bm{x}_{s}^{\top})\in\mathcal{R}^{sd}$ has $s$ tokens with token feature $\bm{x}_{i}\in\mathcal{R}^{d},\forall i\in[s]$, and label $y\in\mathcal{N}^{+}$. The objective is to learn the map of $\bm{x}$ to the corresponding $y$. The general MoE structure are formulated as
@@
where $n$ is the number of experts, $G(\bm{x}_{t})\colon\mathcal{R}^{d}\to\mathcal{R}^{n}$ is the gating weight vector of experts which maps the tokens of $\bm{x}_{t}$ into the coresponding experts with weights, e.g., $G_{i}(\bm{x})=\text{Softmax}(\bm{W}\bm{x}+\bm{\epsilon})$ where the softmax is applied to each row, and $E_{i}(\bm{x}_{t})\colon\mathcal{R}^{d}\to\mathcal{R}$ is the $i$-th expert network, see (Liu et al. 2024) for current different router methods. Generally, $n\ll s$, which saves much computation compared to the dense structure.
Expert-Token Affinity.
$\bm{W}_{\text{aff}}$ denotes the expert-token affinity matrix. With GrAP as the layer of feature extraction, the formulation of $\bm{W}_{\text{aff}}$ is as followed:
$\bm{W}_{\text{aff}}=\begin{pmatrix}\bm{w}_{1}&0&\cdots&0\\ 0&\bm{w}_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\bm{w}_{n}\\ \end{pmatrix}$ (2)
$\bm{w}_{i}=\frac{n}{d}\cdot\mathbf{1}\left\{\frac{i\cdot d}{n}\leq j<\left(i+1\right)\frac{d}{n}\right\}\quad 0\leq j<d$ (3)
The expert-token affinity matrix is employed as the gating weight to calculate the affinity score between each expert and token. We define the affinity score of $t$-th token and $i$-th expert as the cosine similarity between vectors $\bm{x}_{t}$ and $\bm{w}_{i}$:
$\delta_{ti}=\cos\left(\bm{x}_{t},\bm{w}_{i}\right):=\bm{x}_{t}^{\top}\bm{w}_{i}/(\|\bm{x}_{t}\|\cdot\|\bm{w}_{i}\|)$ (4)
The affinity score intuitively reflect how closely the two inputs are associated. From a perspective of semantic, the affinity scores derived from affinity metrics consisting of orthogonal vectors represent the degree of association between each token and various experts, as shown in Figure 1. Therefore, we leverage the affinity score as the principle of our affinity-driven active selection routing mechanism.
Figure 1: The illustration of affinity score.
Figure 2: The architecture of the gate network along with the hybrid TCR + ECR router.
@@
We consider our affinity-driven active selection routing as a hybrid of TCR (Clark et al. 2022; Zhou et al. 2022) and ECR. As the name suggested, TCR lets each token choose its top-scored experts, and ECR lets each expert choose its top-scored tokens. Specifically, we use the result of the expert-token affinity metrics as the affinity score between tokens and experts. In conventional TCR routing strategy, the tokens are simply route to their Top-1 expert. In our hybird TCR+ECR routing strategy, experts also select tokens for processing from assigned tokens according to affinity scores:
$\left(\tilde{E}_{t1},\dots,\tilde{E}_{t\ell}\right)=\text{Top-}\ell\left(\left\{\delta_{t1},\dots,\delta_{tn}\right\}\right),$ (5)
$~{}~{}\tilde{I}_{tk}\in[n],\forall t\in[s],k\in[\ell].$
and then the expert to choose its $-\ell$ tokens where $\ell$ is determined by a threshold of the sum of affinity scores:
$\left(I_{1i},\dots,I_{Ci}\right)=\text{Bottom-}C\left(\left\{t\in[s]:\exists j\in[\ell],\tilde{I}_{tj}=i\right\}\right),$ (6)
$~{}~{}I_{ki}\in[s]\cup\text{None},\forall i\in[n],k\in[C].$
Such bidirectional selection mechanism motivates each expert to receive a certain number of tokens with the highest affinity score to itself, thereby achieving a resonance effect. The resonance effect can help mitigate the homogenization in MoE.
Locality Loss.
@@
Feed-forward network (FFN) layers are commonly employed in expert networks, allowing each expert to learn independently as a separate neural network, thus preventing interference between samples. This mechanism leads to a severe load imbalance, as experts frequently selected in the early stages are more likely to be chosen in later stages. To mitigate this skewness in token allocation, the auxiliary loss (Shazeer et al. 2017) has been proposed. Building upon the auxiliary loss, LocMoE introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$. This loss term serves as a soft constraint, encouraging tokens to be sent to experts residing on the same node, thereby mitigating the substantial overhead incurred by partial inter-node communication.
Training Strategy
The Dynamic Token Distribution Evolution.
@@
The hybrid TCR+ECR bidirectional selection routing, introduced in the model structure, is exemplified in the figure. If the feature fragment corresponding to the $k$-th dimension of the gating weight for a particular token is more prominent, then that token will be routed to the $k$-th expert. If among all tokens routed to the $k$-th expert, there is a certain probability of the presence of class-discriminative tokens, then the capacity $C$ must be set to a larger value to ensure the inclusion of sufficient class-discriminative tokens. The router proposed in this paper is a hybrid of TCR and ECR modes. After determining the expert to which a token will be routed, scores are calculated for the tokens assigned to each expert, and a $-\ell$ selection is performed, where $\ell^{*}$ is determined by a threshold of the sum of scores. Subsequent theoretical analysis will demonstrate the effectiveness of this hybrid routing scheme.
Theoretical Explanation
@@
To explain the motivation of our method, we show some theoretical insights in this section. Our theoretical analyis is bulit on Chowdhury et al. (2023), where they make the following data assumption:
Assumption 1 (data assumption).
Each input $\bm{x}\in\mathcal{R}^{sd}$ with $s$ tokens is comprised of one class-discriminative pattern $\bm{o}_{1},\dots,\bm{o}_{n}\in\mathcal{R}^{d}$, with each decides the label in $[n]$, and $s-1$ class-irrelevant patterns $\bm{r}\sim\mathcal{N}$ for certain distribution $\mathcal{N}$. For example, $\bm{x}=(\bm{r}_{1},\bm{r}_{2},\bm{o}_{1},\bm{r}_{3},\dots,\bm{r}_{s-1})$ has label $1$, where $\bm{r}_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N},\forall i\in[s-1]$.
Based on Assumption 1, Chowdhury et al. (2023) demonstrated that the training of MoE go through two phases:
@@
Phase 1: Router training (Chowdhury et al. 2023, Lemma 4.1 and Assumption 4.4), which makes class-discriminative patterns all to the corresponding expert. This process ensures that each expert only receives the class-discriminative tokens related to the specific class.
@@
Phase 2: Expert training (Chowdhury et al. 2023, Theorem 4.2 and Theorme 4.5), which makes each expert learn to predict the label based on its class-discriminative inputs from Phase 1. This process is designed to establish each expert’s ability to handle and solve problems.